OpenStudy (anonymous):

HELP!!! > At the beginning of each school year, Madison Middle School orders new shirts for the entire student body. Each student receives one shirt. Each shirt is the same price, s, and the school orders enough for n students. Which equation represents the proportional relationship between P, the total price of the shirts, and the number of students?

OpenStudy (smored):

since the total number (P) is the number of students (n) times the price of shirts (s) the equation would be P=ns

OpenStudy (smored):

well there is one option which says exactly what I said

OpenStudy (anonymous):

lmaoo

OpenStudy (smored):

ya find it?

OpenStudy (anonymous):

A lol

OpenStudy (smored):

thats the answer

OpenStudy (anonymous):

yes sn=P

OpenStudy (anonymous):

Which equation could be used to find the number of minutes, m, in h hours?

OpenStudy (smored):

m=60h since 60min=1hr 1min=1/60hr

OpenStudy (smored):

thank you for them medals.

OpenStudy (anonymous):

np ,

OpenStudy (anonymous):

ok 2 more

OpenStudy (anonymous):

A class is going on a field trip to see a movie. Each ticket is $10.00. The number of students, n, who will receive a $10.00 ticket is proportional to the total cost, C, of all the tickets. Which equation could be used to find the total cost of all the tickets?

OpenStudy (anonymous):

@smored

OpenStudy (smored):

each n gets $10.00 and that equals C C=n$10.00

OpenStudy (anonymous):

Does the table below represent a proportional relationship? x 2 3 4 5 y 6 7 8 9

OpenStudy (anonymous):

No, because all the ratios of y to x are not equal. Yes, because all the ratios of y to x are equal to 3. Yes, because all the ratios of y to x are equal to 2. No, because all the numbers are positive.

OpenStudy (smored):

the first one. 6/2=3 7/3=2.333 8/4=2 9/5=1.8

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Does the table below represent a proportional relationship? x 2 3 4 5 y -8 -12 -16 -20 No, because the ratios of y to x are not equal. Yes, because all the ratios of y to x are equal to -4. No, because the table has negative numbers. Yes, because all the ratios of y to x are equal to -2.

OpenStudy (smored):

2nd answer (divide y by x and you will always get -4)